HWill sec 6.4 Previous 1 2 3 4 5 6 Next Question 4 of 6 (1 point) 64 Section Exercise9 I In a recent year, about 22% of Americans 15 years and older are single, what is the probability that in a random sample of 240 Americans 15 or older, more than 36 are single? Round the final answer to 4 decimal places and intermediate z-value calculations to 2 decimal places. P(X > 36-10-9817 As Li Sa

this is wrong answer
Answer #1

Given p = 0.22 n = 240

μ = n * p = 240 * 0.22 = 52.8

σ = √(np(1-p)) = √(52.8 * 0.78) = 6.4175

P(X > 36) = P(z > (36 - 52.8) / 6.4175)

= P(z > -2.62)

Using z tables, the value is 0.9956.